1.000 mole of H2 gas and 1.000 mole of I2 vapor are introduced into a 5.00-liter sealed flask. The mixture is heated to a certain temperature and the following reaction occurs until equilibrium is established.
H2(g) + I2(g) ⇄ 2HI(g) 2-1.58=0.42/2=0.21
At equilibrium, the mixture is found to contain 1.580 mole of HI. (a) What are the concentrations of H2, I2 and HI at equilibrium? (b) Calculate the equilibrium constant Kc.
1
Expert's answer
2020-11-26T05:09:43-0500

H2(g)+I2(g)=2Hl(g)

I (moles)        1        1        0

E(moles)       1-x    1-x       2x

                                1.580

Moles of H2 that reacted= 1.580/2

                                      =0.79 moles

Moles of I2 that reacted =0.79 moles

Moles of H2 at equilibrium= 1-0.79

                                          =0.21 moles

Moles of I2 at equilibrium= 0.21 moles

Concentration of H2 at equilibrium= 0.21/5

                                                      =0.042moldm-3

Concentration of I2 at equilibrium= 0.042moldm-3

Concentration of HI at equilibrium=1.580/5

                                                       =0.316 moldm-3

Kc= [HI]2/{[H2] [I2]}

   = (0.316)2/ {0.042×0.042}

   = 56.61